Penn State shield
Skip to content Skip to search

Headlines

ITS - New Kensington Alerts

/alerts March 29 @ 5:19 am

ITS Alerts by Date


ITS Alerts by Service

ITS Alerts by Location

  • red boxCurrent Alert
  • green boxResolved Alert
  • orange boxFuture Alert

New Kensington RSS badge

The most recent ITS Alerts for New Kensington are listed below and are also published via the New Kensington Alerts RSS feed. In addition to the alerts below, a continuously updated list of the latest virus threats is provided on the right side of this page.

red boxANGEL Logins Degraded

Last updated on March 28, 2015 at 10:29:45 PM

At 7:59 on March 28, 2015 ANGEL experienced a degradation in service for users who attempted to log in. Users already in ANGEL were unaffected. ITS staff have changed the...read more

green boxResolved: Meeting @ PennState temporarily unavailable this morning

Last updated on March 27, 2015 at 08:47:29 AM

This morning, from approximately 7:30am to 8:08am the Meeting @ PennState service was unavailable. Staff are researching the outage and will update this ticket as soon as there is more...read more

green boxResolved: 800/Toll Free Numbers

Last updated on March 25, 2015 at 11:02:18 AM

ITS/TNS is experiencing issues with our 800/Toll Free provider. Currently all numbers are pointing to the incorrect numbers. We are working with our provider to get the problem resolved in...read more

green boxResolved: Resolved: Internet service disruption.

Last updated on March 25, 2015 at 08:40:33 AM

An unplanned circuit vendor outage this morning, impacted the University's Internet connectivity between 1:03 AM and 2:37 AM. ITS is continuing to investigate the nature of this outage with our...read more

orange boxDataFlux Maintenance

Last updated on March 24, 2015 at 03:36:55 PM

On Tuesday, March 31, 2015, during the regularly scheduled maintenance window (5:00 a.m. - 7:00 a.m.), Information Technology Services (ITS) will perform maintenance on the DataFlux production server. Services using...read more

green boxResolved: University Collaborative Suite (UCS) Partial Outage

Last updated on March 24, 2015 at 03:25:36 PM

From 2:14-2:29 p.m., UCS servers experienced a partial outage. Staff is looking into the root cause of the issue. No e-mail was lost as a result of the outage.

green boxResolved: DataFlux Service Changes

Last updated on March 24, 2015 at 08:46:56 AM

On Monday March 23, 2015, at 7 a.m., all AIS DataFlux match code requests will no longer be served by the legacy AIS DataFlux infrastructure. At this time, all requests...read more

green boxResolved: IBIS/ISIS Disaster Recovery Exercise - March 11-20, 2015

Last updated on March 20, 2015 at 09:05:35 PM

ITS successfully performed this disaster recovery exercise. By 6:00 a.m. this morning IBIS and ISIS were returned to their individual enterprise servers with verification testing completed by 7:00 a.m.

green boxResolved: Internet service degradation

Last updated on March 19, 2015 at 12:00:53 PM

This maintenance has been completed. Internet service returned to normal operations at 1:00 AM.

green boxcactus.aset.psu.edu Maintenance

Last updated on March 18, 2015 at 03:19:36 PM

On Thursday, March 19, during the maintenance window (5-7 a.m.), an upgrade will be made to the cactus.aset.psu.edu server. A short, minute-long outage may occur while maintenance is being performed....read more

More ->